# IOVS4NeRF: Incremental Optimal View Selection for Large-Scale NeRFs

### v1

@@ IV-B Hybrid-Uncertainty Estimation
Therefore, our rendering uncertainty can be expressed as:
$\bullet$ Positional Uncertainty.
In deep learning, uncertainty estimation methods mainly include Bayesian estimation[13] and model ensemble[14]. We found that the basic NeRF ensemble should not only calculate the variance in the volume rendering space but also consider an additional positional information term. We observed that the trajectories of most drone photography are planar[41]. To improve our generalization performance, we also incorporate non-planar trajectories into the measurement.
For planar trajectories, we use the Voronoi information gain radiation field, which is a hierarchical planner consisting of a top-level planner that forms local path points using Voronoi vertices[42] and a bottom-level planner that refines uncertainty biases and transfers them to the top level for selection, as shown in Fig. 4.
The top-level global planner uses an improved version of the generalized Voronoi diagram to form a graph of collision-free space. By traversing the nodes in the graph, the global planner creates and updates a list of Voronoi nodes, forming pose points that cover the maximum uncertainty. Then, the bottom-level local planner selects the largest node from the list based on the area potential field formed by weighted centroids and uploads it to the top-level planner for further filtering. This allows for adaptive exploration of position uncertainty. Specifically, we compress three-dimensional pose information points to two-dimensional Euclidean space and define ${p_{i}}$ as n distinct points on this space, and $\lambda_{i}$ as the weighted value of a given point. Then, ${{\mathop{\rm V}\nolimits}_{n}}({p_{i}},{\lambda_{i}})$ is the V-region of point ${p_{i}}$ with weight $\lambda_{i}$ , where ${d(p,{p_{i}})}$ is the Euclidean distance between ${p}$ and ${p_{i}}$:
@@
${F_{p}}(I)=\sum\limits_{i=1}^{{N_{v}}}{\frac{{\sum\limits_{j=1}^{{N_{v}}}{||{p_{i}}-{p_{j}}|{|^{{\lambda_{i}}}}}}}{{{A_{i}}}}}$ (13)
, where ${{A_{i}}}$ represents area of the Voronoi polygon ${{V_{n}}({p_{i}})}$, and ${{N_{v}}}$ means the total number of 3D pose points.
For non-planar flight trajectories, we employ the Voronoi clustering algorithm[43], which provides a threshold as an input parameter. This threshold represents the maximum volume allowed for units that can still be combined into evolving clusters. We approximate local instance space density through unit volume, making it important only when determining if the density is high enough to further combine into clusters, as shown in Fig. 5.
@@
The purpose of point cluster generalization is to correctly convey large-scale information into small-scale information diagrams, considering topological information, and metric information[43] in the process. Based on this, we combine the concepts of importance value and relative local density with the aforementioned cell cluster theory to evaluate the change in the importance value of the entire area and the importance value of the variation. The larger the area of the Voronoi polygon, the greater the weight. Thus, points are more likely to be retained in generalized mapping. The importance value equation is as followed:
${G_{i}}=\frac{{{\lambda_{i}}{A_{i}}}}{{\sum\limits_{k=1}^{{N_{v}}}{({\lambda_{k}}{A_{k}})}}}$ (14)
The probability of selecting the specific point, denoted as $G_{i}$, is given by the product of the area of the Voronoi polygon, denoted as $A_{i}$, and the weight value, denoted as $\lambda_{i}$.
Relative local density enables the comparison of density changes point by point before and after generalization, thus better assessing the density changes between points before and after generalization.
@@
The position information uncertainty of non-planer trajectory is:
@@
The positional uncertainty can be expressed as:
$\sigma_{pos}^{2}(\textsl{I})=\mathbb{I}(p){F_{p}}(\textsl{I})+\mathbb{I}(np){F_{np}}(\textsl{I})$ (17)
and $\mathbb{I}(\cdot)$ represents the indicator function.
@@ IV-C Implementation
In this section, we will shed some insight on our project. The first introduction is that we propose proportionally distributing samples for improved rendering effects while accelerating sampling based on a threshold. Specifically, we accelerate the process by accumulating transmittance along the ray and terminating sampling beyond a certain threshold.
We also design a classifier to determine whether the trajactory is planar or non. To implement our adaptive classifier, we divide the data point set into two groups, $A={a_{1},...,a_{p}}$ and $B={b_{1},...,b_{q}}$, and measure the similarity between these two sets using the Hausdorff distance. Given two trajectories $A=a_{i}$ and $B=b_{j}$, the Hausdorff distance between sets $A$ and $B$ is defined as:
$H(A,B)=\max(h(A,B),h(B,A))$ (18)
, where
@@
Compared to traditional distance metrics, the Hausdorff formula allows direct computation of trajectory similarity without requiring interpolation of trajectory sets, avoiding the addition of noise to trajectory data. By setting a fixed threshold, we can determine the flight trajectory. If the Hausdorff distance between sets $A$ and $B$ exceeds the threshold, indicating low similarity, it can be classified as a non-planar trajectory; otherwise, it is a planar trajectory.
1) Initialization for poses: We input the ground truth RGB images $\{I_{i=1}^{N}\}$ into COLMAP, which is a software to recover the structures from images and output 3D position $\mathbf{x}=(x,y,z)$ in the scene and a viewing direction vector $\mathbf{d}=(\theta,\phi)$.
2) Initialization for NeRF: We randomly select a specific amount of images from dataset to train our weights of NeRF MLP network. Through experiments, we recommend initializing with a quantity that is 10 percent of the total, and an absolute minimum of 20 images.
3) The usage of the variants of NeRFs in our experiments: We use CF-NeRF. ActiveNeRF and D.A.Ensemble as baseline comparisions. In these, ActiveNeRF and CF-NeRF are hard-coded, which change the internal parameters and the uncertainty of the neural network’s output. Thus it cannot be directly embedded into our IOVS framework. However, modifying some of the outputs still allows for integration just as D.A.Ensemble achieves.
@@
Algorithm 1 Hybrid-Uncertainty Calculation Algorithm
@@
2:Output: hybrid-uncertainty
3:Compute 3D position $\mathbf{x}$, and viewing direction $\mathbf{d}$ using COLMAP
@@
6:Construct Beta Distribution based on Eq. (11) and output $\bar{C}(\textbf{r})$, $\bar{\beta}^{2}(\textbf{r})$
@@
18:Normalize and compute hybrid-uncertainty based on Eq. (6)
@@ V-A Experimental Setup
Datasets. Our dataset comprises seven groups: two benchmarks from Mill 19, two from Urban 3D, one from Pix-4D, and two self-captured footage datasets from Changsha, Hunan.
We compare all the methods in the following publicly available large-scale datasets caputred by UAVs:
@@
  • Villa: villa is a UAV photogrametry dataset collected and publiced by Pixel4D, which includes 232 images with the resolution at 6000$\times$4000 pixels  [^3] https://earldudley.com/products/pix4dmatic/.
@@
Metrics. Note that this work focuses on optimal view selection with hybrid-uncertainty estimation. Thus, our evaluation metrics are divided into two types.
@@
Comparison methods. We compare the proposed solution IOVS4NeRF over three state-of-the-art large-scale NeRF solutions with a dynamic view selection strategy to leverage the reconstruction quality and computational consuming:
@@
Setup Details. We utilize the Instant-NGP [17] to demonstrate the effectiveness of the proposed IOVS4NeRF. But please note that as we illustrate in section XXX we can replace Instant-NGP with any NeRF solutions (supplemental material). All the experiments are conducted with a Intel core I9 CPU and an NVIDIA GeForce 3090 GPU (24GB memory).
TABLE I: Quantitative results in 6 real image scens: IOVS4NeRF performs superior to or on par with the three state-of-the-art related solutions CF-NeRF, Active-NeRF and D.A.Ensenble in the quality of synthesized novel views (PSNR, SSIM and LPIPS), the accuracy of uncertainty estimation (AUSE and SRCC) and the efficiency (time-consuming). The best and second results are shown in green and blue, respectively.
@@ V-B Evaluation of Uncertainty Estimation
We compare IOVS4NeRF against three SOTA NeRF-Based view selection solutions. As shown in Table. I, IOVS4NeRF achieves significant more information in uncertainty prediction with respect to synthesis error over CF-NeRF, Active-NeRF and D.A.Ensemble. The superior performance of our the approach demonstrates the proposed hybrid-uncertainty leads to more consistent uncertainty estimates compared to solely rendering-based uncertainty such as CF-NeRF, Active-NeRF, and D.A.Ensemble.
@@ V-C Evaluation of Novel View Synthesis
For each dataset, due to GPU limitations, we randomly select 500 images as the full image set (if the total number of image in a dataset is less than 500, we use all the images instead). We randomly choose 15% of the images in each dataset for initialization and randomly select 10% of the images as the test set. Then we incrementally select 15% of the images using vrious view selection methods as the optimal views for incremental training. Thus, only 30% of the full image set are used to synthesis the novel view, of which quality are evaluted by using PSNR, SSIM and LPIPS.
@@
Figure 7: Comparison of IOVS4NeRF against to two baselined view selection strategies random and FVS on the ArtSci dataset. For each test scene, we incrementally re-train the instan-NGP model by adding the new views selected by IOVS4NeRF, random and FVS at each step, then render tested views and report the average PSNR scores with standard deviations over all test scenes and runs respectively. Note that the large standard deviations are due to the varying rendering difficulty of each scene. Our proposed approach IOVS4NeRF finds informative images in the scene, improving scene representations via image-based neural rendering.
Figure 8: Qualitative comparisons in ablation study by removing the rendering uncertainty $\sigma_{rgb}^{2}$ and positional uncertainty $\sigma_{pos}^{2}$ in proposed hybrid-uncertainty one by one.
(1) Comparison of Baseline Optimal View Selection Strategies: We compare the proposed optimal view selection approach IOV4NeRF against two heuristic baselines:
  • Random: selects a view candidate uniformly at random.
  • FVS: selects the view that maximizes the view distance with respect to previously collected images.
@@
(2) Comparison of State-of-the-art Optimal View Selection: We validate the performance of our proposed framework, IOVS4NeRF, and compare it with three state-of-the-art NeRF-based optimal view selection solutions, namely, CF-NeRF, Active-NeRF and D.A.Ensemble.
We first show the results with an incremental learning scheme, where the time and computation resources are considered sufficient. As shown in Figure. 6, we can easily see that IOVS4NeRF generate better visual quality synthesized novel views against the outputs of CF-NeRF, Active-NeRF and D.A.Ensemble, that are closed to the ground truth among all 6 datesets. The quantitative results in Table. I also show IOVS4NeRF achieves the best performance in three synthesis image quality evaluation metrics (PSNR, SSIM, LPIPS) over the CF-NeRF, Active-NeRF and D.A.Ensemble. These experimental observations prove the proposed method IOVS4NeRF can select most informative inputs compared with state-of-the-art approaches, which contributes most to synthesizing views from less observed regions.
The runtime comparison in Table. I shows that IOVS4NeRF outperforms CF-NeRF, Active-NeRF and D.A.Ensemble in efficiency with a significant margin. We interpret this acceleration are mainly due to all three compared optimal view selection strategies in CF-NeRF, Active-NeRF and D.A.Ensemble are hard-coded with a NeRF module while IOVS4NeRF is a flexible framework where the proposed hybrid-uncertainty computation is as a plug-in to any NeRF solution (e.g. Instant-NGP in our experiment). This soft-coded scheme leads to an advantage in efficiency. It is vital to notice that once a more advanced NeRF solution be developed in the future, it can be easily applied in IOVS4NeRF and is supposed to achieve further improvements in both synthesize quality and time-consuming.
TABLE III: Ablation on hybrid-uncertainty by removing the rendering uncertainty $\sigma_{rgb}^{2}$ and positional uncertainty $\sigma_{pos}^{2}$ one by one. The best and second results are shown in green and blue, respectively.
Method Scene Building Rubble Polytech ArtSci Villa University
@@
V-D Ablations Study
Note that in this paper, we propose to use hybrid-uncertainty over the solely rendering uncertainty that is widely used in existing works. Thus, in this experiment, we further evaluate the effectiveness of the candidate uncertainty estimation module alone.
We denote the baseline approaches as:

### v2

@@ IV-B Hybrid Uncertainty Estimation
Assuming the distribution of the presented rays is independent, we can measure the information content of the viewpoint image by calculating the maximum log-likelihood of the rays from the same image. Therefore, our rendering uncertainty can be expressed as:
$\sigma_{rgb}^{2}(I)=\sum\limits_{i=1}^{{N_{r}}}{\frac{{||C({{\bf{r}}_{i}})-\bar{C}({{\bf{r}}_{i}})||_{2}^{2}}}{{2{{\bar{\beta}}^{2}}({{\bf{r}}_{i}})}}+\frac{{\log{{\bar{\beta}}^{2}}({{\bf{r}}_{i}})}}{2}},$ (12)
where $N_{r}$ means the total number of the rays in photo I.
Positional Uncertainty. It has been determined that most drone flight paths that can be programmed follow an approximate planar trajectory, but users can freely manipulate the drone to fly along non planar trajectories. Therefore, it is necessary to consider the position distribution of drone trajectories.itemize
$\bullet$ Planar Trajectories. We employ a Voronoi-based information gain radiance field within a hierarchical planning framework. This framework comprises a superior planner that delineates local trajectory points utilizing Voronoi vertices, as per the methodology outlined in [32]. Concurrently, an inferior planner is tasked with the refinement of uncertainty biases, subsequently elevating these for consideration by the superior planner. The schematic representation of this process is depicted in Figure 4.
The top-level global planner uses an improved version of the generalized Voronoi diagram to form a graph of collision-free space. By traversing the nodes in the graph, the global planner creates and updates a list of Voronoi nodes, forming pose points that cover the maximum uncertainty. Then, the bottom-level local planner selects the largest node from the list based on the area potential field formed by weighted centroids and uploads it to the top-level planner for further filtering. This allows for adaptive exploration of uncertainty. Specifically, we compress 3D pose information points to two-dimensional Euclidean space and define ${p_{i}}$ as n distinct points on this space, and $\lambda_{i}$ as the weighted value of a given point. Then, ${{\mathop{\rm V}\nolimits}_{n}}({p_{i}},{\lambda_{i}})$ is the V-region of point ${p_{i}}$ with weight $\lambda_{i}$ , where ${d(p,{p_{i}})}$ is the Euclidean distance between ${p}$ and ${p_{i}}$:
@@
${F_{p}}(I)=\sum\limits_{i=1}^{{N_{v}}}{\frac{{\sum\limits_{j=1}^{{N_{v}}}{||{p_{i}}-{p_{j}}|{|^{{\lambda_{i}}}}}}}{{{A_{i}}}}},$ (14)
where ${{A_{i}}}$ represents area of the Voronoi polygon ${{V_{n}}({p_{i}})}$, and ${{N_{v}}}$ means the total number of 3D pose points.
$\bullet$ Non-planar Flight Trajectories. We employ the Voronoi clustering algorithm[33], which provides a threshold as an input parameter. This threshold represents the maximum volume allowed for units that can still be combined into evolving clusters. We approximate local instance space density through unit volume, making it important only when determining if the density is high enough to further combine into clusters, as shown in Fig. 5.
@@
The purpose of point cluster generalization is to correctly convey large-scale information into small-scale information diagrams, considering topological information, and metric information in the process. Based on this, we combine the concepts of importance value and relative local density with the aforementioned cell cluster theory to evaluate the change in the importance value of the entire area and the importance value of the variation. The larger the area of the Voronoi polygon, the greater the weight. Thus, points are more likely to be retained in generalized mapping. The importance value equation is as followed:
${G_{i}}=\frac{{{\lambda_{i}}{A_{i}}}}{{\sum\limits_{k=1}^{{N_{v}}}{({\lambda_{k}}{A_{k}})}}},$ (15)
where $G_{i}$ is the probability of selecting the specific point and given by the product of the area of the Voronoi polygon, denoted as $A_{i}$, and the weight value, denoted as $\lambda_{i}$.
Relative local density enables the comparison of density changes point by point before and after generalization, thus better assessing the density changes between points before and after generalization. As follows:
@@
The information uncertainty of non-planer trajectory is:
@@
The all uncertainty can be expressed as:
$\sigma_{pos}^{2}(\textsl{I})=\mathbb{I}(p){F_{p}}(\textsl{I})+\mathbb{I}(np){F_{np}}(\textsl{I}),$ (18)
where $\mathbb{I}(\cdot)$ represents the indicator function.
@@ IV-C Implementation
In this section, we introduce the implementation details of the proposed IOVS4NeRF.
Proportionally Distributing Samples. The first introduction is that we propose proportionally distributing samples for improved rendering effects while accelerating sampling based on a threshold. Specifically, we accelerate the process by accumulating transmittance along the ray and terminating sampling beyond a certain threshold.
Trajectory Strategy Classifier. We also design a classifier to determine whether the trajectory is planar or non. To implement our adaptive classifier, we divide the data point set into two groups, $A={a_{1},...,a_{p}}$ and $B={b_{1},...,b_{q}}$, and measure the similarity between these two sets using the Hausdorff distance. Given two trajectories $A=a_{i}$ and $B=b_{j}$, the Hausdorff distance between sets $A$ and $B$ is defined as:
$H(A,B)=\max(h(A,B),h(B,A)),$ (19)
where,
@@
Compared to traditional distance measurement methods, the Hausdorff distance formula can directly calculate trajectory similarity without the need to interpolate the trajectory set, thereby avoiding the introduction of noise. By setting a fixed threshold, the nature of the flight trajectory can be determined. If the Hausdorff distance between sets $A$ and $B$ exceeds the threshold, it indicates low similarity and the trajectory can be classified as non-planar; otherwise, it is considered planar.
Solution Implementation. As shown in Alg. 1 that the implementation of proposed IOVS4NeRF has four main steps:
1) Initialization for poses: We input the ground truth RGB images $\{I_{i=1}^{N}\}$ into COLMAP, which is a software to recover the structures from images and output 3D $\mathbf{x}=(x,y,z)$ in the scene and a viewing direction vector $\mathbf{d}=(\theta,\phi)$.
2) Initialization for NeRF: We randomly select a specific amount of images from dataset to train our weights of NeRF’MLP network. Through experiments, we recommend initializing with a quantity that is 15% of the total.
3) The usage of the variants of NeRFs: Chosing a lightweight NeRF model can further accelerate the reconstruction speed. Specifically, we apply Instant-NGP to the framework in our experiments.
@@
Algorithm 1 Hybrid Uncertainty Calculation Algorithm
@@
2:Output: hybrid uncertainty
3:Compute 3D $\mathbf{x}$, and viewing direction $\mathbf{d}$ using COLMAP
@@
6:Construct Gaussian Distribution based on Eq. (9) and output $\bar{C}(\textbf{r})$, $\bar{\beta}^{2}(\textbf{r})$ based on Eq. (10), Eq. (11)
@@
18:Normalize and compute hybrid uncertainty based on Eq. (7)
@@ V-A Experimental Setup
Datasets. Our dataset comprises 7 groups: 2 from Mill-19, 2 from Urban-3D, one from Pix-4D, and 2 self-captured footage datasets from Changsha, Hunan.
We compare all the methods in the following publicly available large-scale datasets captured by UAVs:
@@
  • Villa: Villa is a UAV photogrametry dataset collected and released by Pixel4D, which includes 232 images with the resolution at 6000$\times$4000 pixels  [^3] https://earldudley.com/products/pix4dmatic/.
@@
Metrics. Note that this work focuses on optimal view selection with hybrid uncertainty estimation. Thus, our evaluation metrics are divided into two types.
@@
Comparison Methods. We applied the proposed solution, IOVS4NeRF, and three current popular methods designed to exclusively estimate uncertainty in NeRF-based models to large-scale scenes for a comparison of reconstruction quality and computational consuming:
@@
Setup Details. We utilize the Instant-NGP [9] to demonstrate the effectiveness of the proposed IOVS4NeRF. All the experiments are conducted with a Intel core I9 CPU and an NVIDIA GeForce 3090 GPU (24GB memory).
TABLE I: Quantitative results in 6 real image scenes: IOVS4NeRF performs superior to or on par with the three state-of-the-art related solutions CF-NeRF, Active-NeRF and D.A.Ensemble in the quality of synthesized novel views (PSNR, SSIM and LPIPS), the accuracy of uncertainty estimation (AUSE and SRCC) and the efficiency (time-consuming). The best and second results are shown in green and blue, respectively.
@@ V-B Evaluation of Uncertainty Estimation
We compare IOVS4NeRF against the SOTA of NeRF-Based view selection solutions. As shown in Table. I, IOVS4NeRF achieves significant more information in uncertainty prediction with respect to synthesis error over CF-NeRF, Active-NeRF and D.A.Ensemble. The superior performance of our the approach demonstrates the proposed hybrid uncertainty leads to more consistent uncertainty estimates compared to solely rendering-based uncertainty such as CF-NeRF, Active-NeRF, and D.A.Ensemble.
@@ V-C Evaluation of Novel View Synthesis
For each dataset, due to GPU limitations, we randomly select 500 images as the full image set (if the total number of image in a dataset is less than 500, we use all the images instead). We randomly choose 15% of the images in each dataset for initialization and randomly select 10% of the images as the test set. Then we incrementally select 15% of the images using various view selection methods as the optimal views for incremental training. Thus, only 30% of the full image set are used to synthesis the novel view, of which quality are evaluated by using PSNR, SSIM and LPIPS.
@@
Figure 7: Comparison of IOVS4NeRF against to two baselined view selection strategies Random and FVS on the ArtSci dataset. For each test scene, we incrementally retrain the instant-NGP model by adding the new views selected by IOVS4NeRF, Random and FVS at each step, then render tested views and report the average PSNR scores with standard deviations over all test scenes and runs respectively. Note that the large standard deviations are due to the varying rendering difficulty of each scene. Our proposed approach IOVS4NeRF finds informative images in the scene, improving scene representations via image-based neural rendering.
Figure 8: Qualitative comparisons in ablation study by removing the rendering uncertainty $\sigma_{rgb}^{2}$ and positional uncertainty $\sigma_{pos}^{2}$ in proposed hybrid uncertainty one by one.
Comparison of Baseline Optimal View Selection Strategies. We compare the proposed optimal view selection approach IOV4NeRF against two heuristic baselines:
  • Random: We select a view candidate uniformly at random.
  • FVS: We select the view that maximizes the view distance with respect to previously collected images.
@@
Comparison of State-of-the-art Optimal View Selection. We validate the performance of our proposed framework, IOVS4NeRF, and compare it with three state-of-the-art NeRF-based optimal view selection solutions, namely, CF-NeRF, Active-NeRF and D.A.Ensemble.
We first show the results with an incremental learning scheme, where the time and computation resources are considered sufficient. As shown in Figure. 6, we can easily see that IOVS4NeRF generate better visual quality synthesized novel views against the outputs of CF-NeRF, Active-NeRF and D.A.Ensemble, that are closed to the ground truth among all 6 datasets. The quantitative results in Table. I also show IOVS4NeRF achieves the best performance in three synthesis image quality evaluation metrics (PSNR, SSIM, LPIPS) over the CF-NeRF, Active-NeRF and D.A.Ensemble. These experimental observations prove the proposed method IOVS4NeRF can select most informative inputs compared with state-of-the-art approaches, which contributes most to synthesizing views from less observed regions.
The runtime comparison in Table. I shows that IOVS4NeRF outperforms CF-NeRF, Active-NeRF and D.A.Ensemble in efficiency with a significant margin. We interpret this acceleration mainly due to that all three compared optimal view selection strategies in CF-NeRF, Active-NeRF and D.A.Ensemble are hard-coded with a NeRF module while IOVS4NeRF is a flexible framework where the proposed hybrid uncertainty computation is as a plug-in to any NeRF solution (e.g. Instant-NGP in our experiment). This soft-coded scheme leads to an advantage in efficiency. It is vital to notice that once a more advanced NeRF solution be developed in the future, it can be easily applied in IOVS4NeRF and is supposed to achieve further improvements in both synthesize quality and time-consuming.
TABLE III: Ablation on hybrid uncertainty by removing the rendering uncertainty $\sigma_{rgb}^{2}$ and positional uncertainty $\sigma_{pos}^{2}$ one by one. The best and second results are shown in green and blue, respectively.
Component Scene Building Rubble Polytech ArtSci Villa University
@@
V-D Ablation Study
Note that in this paper, we propose to use hybrid uncertainty over the solely rendering uncertainty that is widely used in existing works. Thus, in this experiment, we further evaluate the effectiveness of the candidate uncertainty estimation module alone. We denote the baseline approaches as: